OpenStudy (gabylovesyou):

Simplify the given expression: (−5 + 3i) • (1 − 2i) −5 − 6i 1 + 13i 2 + i 4 − 5i

OpenStudy (gabylovesyou):

@sleepyjess @SolomonZelman

OpenStudy (gabylovesyou):

1 + 13i ?

OpenStudy (anonymous):

May i help?

OpenStudy (gabylovesyou):

@jhonyy9

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A :)

OpenStudy (anonymous):

I did the expression and it came out as A

OpenStudy (anonymous):

fan and medal please and @ me for more help

OpenStudy (anonymous):

@Gabylovesyou

OpenStudy (solomonzelman):

`\(\large\color{black}{ \displaystyle (-5+3i) \cdot (1-2i) }\) ` \(\large\color{black}{ \Uparrow }\) this piece ` \displaystyle` is just for a proper display in case of a complicated latex. `\(\large\color{black}{ \displaystyle (-5+3i) \cdot (1-2i) }\) ` \(\large\color{black}{ \Uparrow }\) multiplication symbol (dot) the rest I simply wrote. THEN `\(\large\color{black}{ ` `}\)` is where you put your latex ` \large ` is the size ` black ` is the color. all this `\(\large\color{black}{ \displaystyle (-5+3i) \cdot (1-2i) }\) ` results in \(\large\color{black}{ \displaystyle (-5+3i) \cdot (1-2i) }\) `

OpenStudy (solomonzelman):

Sorry for this latex class, but just in case.... it is not hard to do.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle (-5+3i) \cdot (1-2i) }\) re-write it, \(\large\color{red}{ \displaystyle (-5)(1-2i) +(3)(1-2i) }\) then expand and simplify. (beware that \(\large\color{black}{ \displaystyle i^2=-1 }\) )

jhonyy9 (jhonyy9):

@mathmath333 please where is this exercise posted 2-3 days with minute and second hands up of a clock in 24 hours ? - write me please - thank you very much jhonyy9